OpenStudy (anonymous):

GRAPH Project Graphs and Functions Goal: to use equations of lines in a real-world application Role: You work in the marketing department of a tele-communications company Audience: the director of marketing Situation: You are trying to convince the director of marketing to get involved in the cell phone industry. Task: Mobile phone usage has been steadily increasing. The data below shows the mobile phone usage (in millions) base on year: •&νβσπ;1994 – 37.5 million users •&νβσπ;1998 – 70 million users •&νβσπ;2002 – 102.5 million users Performance Task: You are to write

OpenStudy (anonymous):

Please help!!!!!

OpenStudy (anonymous):

I get the linear equation \[y=\frac{65}{8}x+37,5\] where x is equal to time and y is equal to the amount of user at given year.

OpenStudy (anonymous):

Let me try to guide you through this, you have only these three datas given in your problem set. After reading the description I believe that the function is linear, that is a good place to start anyway. So let x be time in years and y be the amount of "million" users. then you can make the following statements about the function: \[x_1 = 0 \ \text{and}\ y_1=q=37.5 \\ x_2=4 \ \text{and} \ y_2= 70\] so you can describe: \[ m = \frac{\Delta y}{\Delta x}= \frac{y_2-y_1}{x_2-x_1} \] and obtain the linear equation \[ y= mx + q \]

OpenStudy (anonymous):

What about the graph and the pridiction ????

OpenStudy (anonymous):

•&νβσπ;1994 – 37.5 million users •&νβσπ;1998 – 70 million users •&νβσπ;2002 – 102.5 million users Performance Task: You are to write a report to the director of marketing that discusses your predications for mobile phone usage and why you believe that it would be good for your company to get involved in the mobile phone industry. Include the following in your report: •&νβσπ;The equation of the line of best for based on the above data •&νβσπ;An interpretation of the value of the slope and the y-intercept •&νβσπ;The reliability of your model •&νβσπ;A prediction for the amount of mobile phone usage two years from today

OpenStudy (anonymous):

the graph is a linear equation, this means you have a linear progression of users. You can obtain the slope with the help I have given you above, therefore you can make a prediction on how steep it is and what that will mean for the users.

OpenStudy (anonymous):

try to find the correct linear equation first, that describes the datas above.

OpenStudy (anonymous):

Then we can discuss it together if you wish.

OpenStudy (anonymous):

What I got was y=32.5/4+37.5

OpenStudy (anonymous):

Where do I go from here??

OpenStudy (anonymous):

Good, so if you want you can graph this function, with a computer or by hand, what you will discover is that the slope is INCREDIBLY steep.

OpenStudy (anonymous):

Ok meaning that the amount of users is increasing

OpenStudy (anonymous):

yes, furthermore it means that it increases A LOT per year. So for your boss/company that would mean he totally wants to invest into that system. Because the profit would be incredible (of course this is only a mathematical model for this exercise though)

OpenStudy (anonymous):

Alright the only other thing I the interpretation of the slope and the y axis

OpenStudy (anonymous):

the slope we have already interpreted

OpenStudy (anonymous):

the y-intercept means x=0

OpenStudy (anonymous):

a steep slope means rapid increase, a flat slope means slow increase

OpenStudy (anonymous):

One last favor can u graph it for me because whenever I do it it comes out weird

OpenStudy (anonymous):

yes okay There you have two identical graphs with different range: http://www.wolframalpha.com/input/?i=%2865%2F8%29x%2B37.5

OpenStudy (anonymous):

Y is it 65/8 instead of 32.5/4?